A small block with mass 0.0400 kg is moving in thexy-plane. The net force on the block is described by the potential- energy function U(x,y)= (5.95 J/m2 )x2-(3.60 J/m3 )y3.

a) What is the magnitude of the acceleration of the block when it is at the point x= 0.30 m , y= 0.56 m?

b)What is the direction of the acceleration of the block when it is at the point x= 0.30 m , y= 0.56 m ?
Answer #1

x componet of force:

F_{x}=-\frac{dU}{dx}=-2(5.95)x

acceleration in x:

F_{x}=-2(5.95)x=ma\rightarrow a=\frac{-2(5.95)0.3}{0.04}

a_{x}=-89.25\frac{m}{s^{2}}

Force in y:

F_{y}=-\frac{dU}{dy}=3(3.60)y^{2}

acceleration in y:

F_{y}=3(3.60)y^{2}=ma\rightarrow a=\frac{3(3.60)0.56^{2}}{0.04}

a_{y}=84.672\frac{m}{s^{2}}

Magnitude of acceleration:

a=\sqrt{89.25^{2}+84.672^{2}}\rightarrow \boldsymbol{a= 123.02\frac{m}{s^{2}}}

------------------------------------------------

direction:

\tan (\theta )=\frac{a_{y}}{a_{x}}=\frac{84.672}{-89.25}

\theta =-43.49^{\circ }   Meassured Clockwise or

\theta =316.51^{\circ } Meassure counterclockwise